OpenStudy (anonymous):

If a transformer steps down a 120-V source to 6.3 V and is connected to an 8 ohm loadWhat is the value of the impedence reflected back to the source?

OpenStudy (anonymous):

How do I figure out the value of impedence reflected back to the source?

OpenStudy (kropot72):

The secondary current is given by \[I _{s}=\frac{6.3}{8}\] The turns ratio n is \[n=\frac{120}{6.3}\] The primary current is given by \[I _{p}=\frac{I _{s}}{n}=\frac{6.3}{8}\times \frac{6.3}{120}\] The impedance reflected back to the source is \[Z _{p}=\frac{V _{p}}{I _{p}}=\frac{120\times120\times8}{6.3^{2}}=8\times(\frac{12}{6.3})^{2}\] So we have the result that impedance reflected back to the source is \[Z _{p}=\frac{R _{s}}{n ^{2}}\]

OpenStudy (kropot72):

\[Z _{p}=8\times(\frac{120}{6.3})^{2}\]

OpenStudy (anonymous):

so 2902 ohms?

OpenStudy (kropot72):

Yes. I get 2902.5 ohms.

OpenStudy (anonymous):

yeah i see how you got that thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

So we have the result that impedance reflected back to the source is \[Z _{p}=R _{l} \times n ^{2}\] where \[R _{l}\ is\ secondary\ load\ resistance\] and \[n ^{2}\ is\ the\ primary/secondary\ turns\ ratio\ squared\]